The standard deviation of the running time of a sample of 200 fuses was found to be



100 hours. Find a 99% confidence interval for the standard deviation of all such fuses.




1
Expert's answer
2021-11-02T10:45:27-0400

The critical values for "\alpha = 0.01" and "df =n-1=200-1=199" degrees of freedom are "\chi_L^2 = \chi^2_{1-\alpha/2,n-1} = 151.3699" and "\chi_U^2 = \chi^2_{\alpha/2,n-1} = 254.1352."

The corresponding confidence interval is computed as shown below:


"CI(Variance)=(\\dfrac{(n-1)s^2}{\\chi^2_{\\alpha\/2,n-1}},\\dfrac{(n-1)s^2}{ \\chi^2_{1-\\alpha\/2,n-1}})\n\u200b"

"=(\\dfrac{(200-1)(100)}{254.1352},\\dfrac{(200-1)(100)}{ 151.3699})\n\u200b"

"=(78.3048, 131.4660)"

Now that we have the limits for the confidence interval, the limits for the 99% confidence interval for the population standard deviation are obtained by simply taking the squared root of the limits of the confidence interval for the variance, so then:


"CI(Standard \ Deviation)"

"=( \sqrt{78.3048}, \sqrt{131.4660})=(8.8490,11.4659)"

Therefore, based on the data provided, the 99% confidence interval for the population variance is "78.3048 < \sigma^2 < 131.4660," and the 99% confidence interval for the population standard deviation is "8.8490 < \sigma < 11.4659."